Can someone help me solve the solution of the system using the elemination method?
0.3x - 0.2y = 4
0.5 + 0.5y = 45/23

Answer by mananth(16946)

0.3x-0.2 y=4 .............1
0.5x+0.5 y=1.96 .............2

multiply (1)by 0.5
Multiply (2) by 0.2
0.15x-0.1y=2
0.1x+0.1 y=0.39
Add the two equations
0.25x=2.39
/0.25
x=9.57
plug value of x in (1)
0.3x-0.2 y=4
2.87-0.2 y=4
-0.2y=4-2.87
-0.2y=1.13
y=-5.65